# Adaptive Selection for Homogeneous Tools: An Instantiation in the RAG Scenario

### v2

@@ Abstract
Current research on tool learning primarily focuses on selecting the most effective tool from a wide array of options, often overlooking cost-effectiveness, a crucial factor in human problem-solving. In this paper, we address the selection of homogeneous tools by predicting both their performance and the associated cost required to accomplish a given task. We then assign queries to the optimal tools in a cost-effective manner. Our experimental results demonstrate that our method achieves higher performance at a lower cost compared to strong baseline approaches.
Adaptive Selection for Homogeneous Tools: An Instantiation in the RAG Scenario
@@ 1 Introduction
Existing tool learning methods focus on selecting the most effective tool from a large number of options Qin et al. (2023b); Tang et al. (2023), while overlooking the crucial aspect of cost-effectiveness, which is a significant criterion in human problem-solving. To bridge this gap, the selection of homogeneous tools has become an important issue. Taking the retrieval-augmented generation (RAG) Lewis et al. (2020); Gao et al. (2023) scenario as an example. Given a user query, a RAG system first uses a search tool to retrieve the background documents. Then, a LLM reads the query and documents to give the response. Given the availability of various web search tools, such as Google and Bing Search, the choice of which tool to use for a given query becomes an important consideration. First, a query that cannot be solved by one tool might be resolved by another. As shown in Figure 1, our dataset reveals that in 4.5% of cases, directly using the LLM to answer queries actually outperforms Google-based RAG. Furthermore, different tools come with varying usage costs. From a cost-effectiveness perspective, it is preferable to select the least expensive tool that can effectively solve the query. If we can adaptively allocate queries to the optimal tool, we can achieve a balance between cost and performance. Additionally, leveraging the complementarity between tools offers the potential for enhanced overall performance. The key challenge in achieving this goal is estimating the performance of candidate tools in handling user queries, instead of making decisions after calling all tools for user queries Chen et al. (2023); Jiang et al. (2023). To overcome this challenge, we automatically construct training data including queries submitted to all tools, along with the returned scores. The training data allows us to learn neural models to predict the performance that LLM calls each tool to solve each query.
@@
Figure 1: The win/tie/lose rates after the LLM uses two compared tools on the TimeQA testset. Besides Bing and Google Search, we consider a non-retrieval condition, and denote the combination as "LLM-only".
@@
In this work, we explore the problem of selecting homogeneous tools and instantiate it in the RAG scenario. We propose a router that assigns input queries to the most appropriate search tool without directly accessing any of the tools. To achieve this, we develop an automated pipeline to construct training data without manual effort. Additionally, we introduce a label-refinement strategy to mitigate noise in labels caused by imperfect information retrieval processes. Finally, we design various assignment strategies to allocate user queries to the optimal tool as needed.
@@
In summary, we make the following contributions. (1) We propose a general approach that dynamically assigns an input query to the optimal tool from a set of homogeneous candidates, only based on the input query. We denote our method as HomoRouter. (2) As far as we know, we are the first to consider the selection of homogeneous tools. Our framework is generalizable to any type of homogeneous tools, extending beyond the RAG scenario. (3) We evaluate our approach on several open-domain QA tasks. Experiments show that our method outperforms the use of a single, fixed search tool across different datasets.
@@ 3.1 Problem Formulation
We demonstrate our work within the RAG framework to provide a realistic illustration. It should be noted that our approach is universal and applicable to various scenarios, not limited to just RAG. Suppose that a LLM needs to solve a set of user queries. The LLM is permitted to call a search tool to generate responses. We consider a situation where there are $M$ homogeneous tools $\{T_{m}\}_{m=1}^{M}$, e.g., Bing and Google Search, each with different costs. Our goal is to develop an adaptive assignment method to achieve a cost-performance trade-off.
@@ 3.2 Training Predicative Model
Figure 2: The framework of our method. Using our predictive model, we predict the scores where the LLM calls each tool to solve each query. Then, we design different strategies to assign each query to the optimal tool on demand.
@@
Label Refinement
It is known that information retrieval is not perfect. The retrieved documents may contain noise, making the labeled score $l_{n}^{m}$ potentially sub-optimal. We adopt the $k$-nearset-neighbors Cover and Hart (1967) strategy to mitigate this issue. For the query $q_{n}$, we use sentence-transformer Reimers and Gurevych (2019) to recall $k$ most similar examples $\{(q_{i},\{l_{i}^{m}\}_{m=1}^{M})\}_{i=1}^{k}$. Then, we obtained the refined label of $q_{n}$ by a linear interpolation between $l_{n}^{m}$ and $\{l_{i}^{m}\}_{i=1}^{k}$:
$\widetilde{l_{n}^{m}}=\alpha\cdot l_{n}^{m}+(1-\alpha)\cdot\frac{1}{k}\sum_{i=1}^{k}l_{i}^{m},\forall m\in\{1,\cdots,M\}.$ (5)
We finally replace $l_{n}^{m}$ (Equation 4) with the refined label $\widetilde{l_{n}^{m}}$ for training the predictive model $\mathcal{M}$.
Methods FreshQA HotpotQA WebQA TimeQA CDQA
$\mathcal{J}(\downarrow)$ Acc. ()($\uparrow$)( ↑ ) $\mathcal{J}(\downarrow)$ Acc. ()($\uparrow$)( ↑ ) $\mathcal{J}(\downarrow)$ Acc. ()($\uparrow$)( ↑ ) $\mathcal{J}(\downarrow)$ Acc. ()($\uparrow$)( ↑ ) $\mathcal{J}(\downarrow)$ Acc. ()($\uparrow$)( ↑ )
Random N/A 83.14 N/A 80.78 N/A 93.23 N/A 68.44 N/A 58.75
LLM-only N/A 76.07 N/A 79.19 N/A 91.14 N/A 50.30 N/A 45.90
LLM+Bing N/A 86.51 N/A 81.76 N/A 93.48 N/A 76.91 N/A 63.86
LLM+Google N/A 86.94 N/A 81.42 N/A 94.97 N/A 77.95 N/A 66.75
HomoRouter (ours) 0.0017 86.49 0.0016 81.35 0.0009 94.68 0.0030 78.49 0.0033 68.19
w/o Refinement (ours) 0.0018 85.05 0.0017 80.99 0.0009 94.42 0.0031 77.69 0.0035 67.89
Table 1: The result of predictive model evaluation. "LLM-only" means the non-retrieval baseline in which the LLM directly responds to the query. "LLM+Bing" and "LLM+Google" means the LLM always call Bing or Google search to solve queries. “Acc. (%)" denotes accuracy. Scores with Bold denote the best values.
@@ 3.3 Assignment Strategies
Given search tools $\{T_{m}\}_{m=1}^{M}$, we denote their usage costs as $\{C_{m}\}_{m=1}^{M}$. [^1] Bing Search charges \$10 for 1000 requests, and Google Search charges \$5 for 1000 requests. Assuming the user has a collection of queries $\{q_{n}\}_{n=1}^{\mathcal{N}}$ that need to be processed, we first use $\mathcal{M}$ to obtain the predicted scores $\{p_{n}^{m}\}_{m=1}^{M}$ for the query $q_{n}$. Then, inspired by Lu et al. (2023); Šakota et al. (2024); Shnitzer et al. (2023), we consider the following strategies to assign queries to corresponding tools.
(1) Fixed-Tool: LLM always calls a fixed tool to handle queries. (2) Best-Predicted: for $q_{n}$ and the predicted $\{p_{n}^{m}\}_{m=1}^{M}$, the LLM selects $T_{m^{\star}}$ $(m^{\star}=\mathop{\arg\max}_{m}p_{n}^{m})$ with the maximal predicted score to handle $q_{n}$. (3) Cost-Saving-ILP: This strategy aims to ensure that the average predicted score, i.e., $p_{n}^{m}$, exceeds a user-defined threshold while minimizing the average cost. This problem can be solved by the integer linear programming (ILP):
@@ 4.1 Experimental Setup
We focus on the open-domain QA scenario. The adopted datasets include public datasets, e.g., HotpotQA Yang et al. (2018), FreshQA Vu et al. (2023), CDQA Xu et al. (2024), WebQA Li et al. (2016) and a private dataset, TimeQA [^2] We leave the detail in Appendix B., that we built ourselves. Data processing details, dataset statistics, and training details are shown in Appendix B.
LLM and Search Tools
We use Qwen-Max [^3] https://dashscope.aliyuncs.com/compatible-mode/v1 Bai et al. (2023) in all of our experiments, due to its strong multilingual ability and lower usage cost. The tools used include Bing and Google Search. In addition, considering that the LLM can answer easy queries independently, we set a non-retrieval condition, i.e., the LLM directly responds to the query. Consequently, its usage cost is set to 0.
@@ 4.2.1 Settings
We also use the Kullback-Leibler (KL) distance of distributions between the predicted $p_{n}^{m}$ and the real $l_{n}^{m}$ as a metric, since a lower distance means a more precise simulation:
$\mathcal{J}=\frac{1}{\mathcal{N}}\sum_{n=1}^{\mathcal{N}}\operatorname{KL}[\text{softmax}_{m}(p_{n}^{m})||\text{softmax}_{m}(l_{n}^{m})].$ (8)
Baselines
We devise following baselines. (1) Random which randomly selects a tool. (2) Fixed-Tool which selects a fixed tool. Since we consider Bing Search, Google Search, as well as a non-retrieval condition, we denote the combinations between the LLM and tools as "LLM-only", "LLM+Bing", "LLM+Google". (3) The ablated variant w/o Refinement that ablates the label-refinement module. It also select the tool with the maximal predicted $p_{n}^{m}$.
@@ 4.2.2 Results
The result is shown in Table 1. Among all the tools, Google Search has the best performance, which coincides with human intuition. "LLM+Bing" and "LLM+Google" outperform LLM-only by a large margin, proving the necessity of utilizing tools for response generation. In addition, we have the following observations.
(1) On FreshQA, HotpotQA and WebQA, our method presents results that are very close to "LLM+Bing" and "LLM+Google", while on TimeQA and CDQA, our method even outperforms the best baseline, i.e., "LLM+Google", with improvements of 0.54% and 1.44%, respectively. This is because despite "LLM+Google" achieving the highest overall accuracy, it still fails to solve some cases that "LLM+Bing" can handle. Specifically, the proportion of such samples is 6.6% in the CDQA testset. Our method can adaptively assign these queries to the optimal tools, leading to better overall results. (2) Our full method, i.e., HomoRouter, outperforms the ablated variant “w/o Refinement", on both metrics of $\mathcal{J}$ and accuracy. This indicates that the label-refinement strategy helps to simulate real scores information.
@@ 4.3 Evaluation of the Adaptively Assignment
According to different strategies in Section 3.3, we assign test queries to the corresponding tools and calculate the average accuracy and usage cost. Cost–accuracy curves are plotted, which present the relationship between the average accuracy and the average cost per query needed.
@@
We plot a cost-accuracy curve on the merged testset of TimeQA and CDQA in Figure 3. "LLM+Google" has a higher accuracy and a lower cost compared to "LLM+Bing". We additionally have the following observations.
(1) Based on our $\mathcal{M}$, "Best-Predicted" achieves a higher accuracy than "LLM+Google", albeit at a higher cost. The accuracy increment comes from the supplement from "LLM+Bing", with a side-effect of higher cost. (2) Compared to both "LLM+Google" and "Best-Predicted", the "Cost-Saving-ILP" strategy is able to achieve a higher accuracy with a lower cost. This shows that "Cost-Saving-ILP" is more flexible and can effectively save cost. More results can be seen in Appendix C.
Figure 3: The cost-accuracy curve on the merged testset.
@@ 4.4 Further Discussion
Impact of $\alpha$ in the Label Refinement Module
We investigate the impact of $\alpha$ (Section 3.2). For the testsets of WebQA, TimeQA, and CDQA, we first use min-max normalization to scale the accuracy of different datasets, then average the corresponding values under different $\alpha$ values. The result is shown in Figure 4. We find that the best result is obtained when $\alpha=0.8$. We hypothesize that $\alpha=0.8$ achieves a good balance between coarse-grained nearest-neighbor-level supervision and fine-grained example-level supervision. We finally set $\alpha=0.8$ for Chinese datasets and $\alpha=0.5$ for English datasets. Similarly, we set $k=2$ for both English and Chinese datasets after searching on the development set.
Figure 4: Impact of $\alpha$ in the label refinement module.
LLM-only LLM+Bing LLM+Google
Raw label $l_{n}^{m}$ 0.71 0.71 0.57
Refined label $\widetilde{l_{n}^{m}}$ 0.62 0.72 0.63
Table 2: Case study for the label refinement.
Case Analysis
We take a case to show why the label-refinement strategy works. For the query from HotpotQA: “Where is the film studio located that produced the 1954 film The Egyptian?", we present its raw labels and refined labels obtained through different tools. As shown in Table 2, due to the noisy retrieval, the raw label obtained from "LLM+Google" are even smaller than "LLM-only". But with the label refinement, the refined label from "LLM+Google" gets larger than that from "LLM-only". In other words, the label refinement allows us to obtain a more robust partial order among multiple homogeneous tools.
@@ 5 Conclusion
Due to the high cost of search tools, our training data only contains approximately 42000 examples, involving three datasets. In the future, with a sufficient budget, we will conduct testing on more datasets. We use Google and Bing search because they are universal search tools that can cover multilingual scenarios. Other search engines, such as Baidu search and Quark search, are not used because they are mainly suitable for the Chinese scenario. We use small encoding models to train our predictive model instead of LLMs, mainly for the following two reasons: (1) Tool selection has a high requirement for latency, while the speed of large models is relatively slow and does not meet the requirement. (2) Tool selection requires the model to predict the score of each tool solving each problem, which is actually a numerical prediction problem. However, the auto-regressive LLMs have significant issues in numerical prediction. In our experiment, the performance of LLMs was inferior to that of small encoding models.
Appendix A Related Work
Tool Learning
LLMs Selection
Appendix B Data Processing Details and Statistics
Construction Details about TimeQA
TimeQA primarily consists of questions related to timeliness. We collected question-answer examples from public QA websites. We prioritize privacy protection and data integrity, ensuring that the dataset contains only open-domain QA content and excludes any private, harmful, or unethical material. TimeQA is used to evaluate LLMs’ capability to answer timeliness questions. Due to confidentiality requirements, we are unable to publicly disclose the dataset.

### v3

# Query Routing for Homogeneous Tools: An Instantiation in the RAG Scenario

@@ Abstract
Current research on tool learning primarily focuses on selecting the most effective tool from a wide array of options, often overlooking cost-effectiveness, a crucial factor in human problem-solving. In this paper, we address query routing for homogeneous tools by predicting both their performance and the associated cost required to accomplish a given task. We then assign queries to the optimal tools in a cost-effective manner. Our experimental results demonstrate that our method achieves higher performance at a lower cost compared to strong baseline approaches.
Query Routing for Homogeneous Tools: An Instantiation in the RAG Scenario
@@ 1 Introduction
Existing tool learning methods focus on routing a query to the most effective tool from a large number of options Qin et al. (2023b); Tang et al. (2023), while overlooking the crucial aspect of cost-effectiveness, which is a significant criterion in human problem-solving. To bridge this gap, query routing for homogeneous tools has become an important issue. Taking the retrieval-augmented generation (RAG) Lewis et al. (2020); Gao et al. (2023) scenario as an example. Given a user query, a RAG system first uses a search tool to retrieve the background documents. Then, a LLM reads the query and documents to give the response. Given the availability of various web search tools, such as Google and Bing, the choice of which tool to use for a given query becomes an important consideration. First, a query that cannot be solved by one tool might be resolved by another. As shown in Figure 1, Our dataset indicates that (1) in 4.5% of cases, directly utilizing the LLM to answer queries surpasses the performance of the Bing-based RAG. And (2) there are 8.1% of instances where the Bing-based RAG fails to deliver a solution, whereas the Google-based RAG successfully provides one. Furthermore, different tools come with varying usage costs. From a cost-effectiveness perspective, it is preferable to select the least expensive tool that can effectively solve the query. If we can adaptively allocate queries to the optimal tool, we can achieve a balance between cost and performance. Additionally, leveraging the complementarity between tools offers the potential for enhanced overall performance. The key challenge in achieving this goal is estimating the performance of candidate tools in handling user queries, instead of making decisions after calling all tools for user queries Chen et al. (2023); Jiang et al. (2023). To overcome this challenge, we automatically construct training data including queries submitted to all tools, along with the returned scores. The training data allows us to learn neural models to predict the performance that LLM calls each tool to solve each query.
@@
Figure 1: The win/tie/lose rates on our PrivateTimeQA test set when the LLM uses two compared tools. Besides Bing and Google, we consider a non-retrieval baseline, and denote the method as "LLM-only".
@@
In this work, we explore query routing for homogeneous tools and instantiate it in the RAG scenario. We propose a router that assigns input queries to the most appropriate search tool without directly accessing any of the tools. To achieve this, we develop an automated pipeline to construct training data without manual effort. Finally, we design various assignment strategies to allocate user queries to the optimal tool as needed.
@@
In summary, we make the following contributions. (1) We propose a general approach that dynamically assigns an input query to the optimal tool from a set of homogeneous candidates, only based on the input query. (2) As far as we know, we are the first to consider the selection of homogeneous tools. Our framework is generalizable to any type of homogeneous tools, extending beyond the RAG scenario. (3) We evaluate our approach on several QA tasks with multiple LLMs. Experiments show that our method outperforms the use of a single, fixed search tool across different datasets.
@@ 3.1 Problem Formulation
We demonstrate our work within the RAG framework to provide a realistic illustration. It should be noted that our approach is universal and applicable to various scenarios, not limited to just RAG. Suppose that a LLM needs to solve a set of user queries. The LLM is permitted to call a search tool to generate responses. We consider a situation where there are $M$ homogeneous tools $\{T_{m}\}_{m=1}^{M}$, e.g., Bing and Google, each with different costs. Our goal is to develop an adaptive assignment method to achieve a cost-performance trade-off.
@@ 3.2 Training Predicative Model
Figure 2: The framework of our method. We first predict the scores where the LLM calls each tool to solve each query. Then, we design different strategies to assign each query to the optimal tool on demand.
@@ 3.3 Assignment Strategies
Given search tools $\{T_{m}\}_{m=1}^{M}$, we denote their usage costs as $\{C_{m}\}_{m=1}^{M}$. Assuming the user has a collection of queries $\{q_{n}\}_{n=1}^{\mathcal{N}}$ that need to be processed, we first use $\mathcal{M}$ to obtain the predicted scores $\{p_{n}^{m}\}_{m=1}^{M}$ for the query $q_{n}$. Then, inspired by Lu et al. (2023); Šakota et al. (2024); Shnitzer et al. (2023), we consider the following strategies to assign queries to corresponding tools.
(1) Fixed-Tool: LLM always calls a fixed tool to handle queries. (2) Best-Performance: for $q_{n}$ and the predicted $\{p_{n}^{m}\}_{m=1}^{M}$, the LLM selects $T_{m^{\star}}$ $(m^{\star}=\mathop{\arg\max}_{m}p_{n}^{m})$ with the maximal predicted score to handle $q_{n}$. (3) Cost-Saving: This strategy aims to ensure that the average predicted score, i.e., $p_{n}^{m}$, exceeds a user-defined threshold while minimizing the average cost. This problem can be solved by the integer linear programming (ILP):
@@ 4.1 Experimental Setup
We focus on the QA scenario. The adopted datasets include public datasets CDQA Xu et al. (2024) and WebQA Li et al. (2016), as well as a private dataset PrivateTimeQA [^1] We leave the detail in Appendix A. that we construct ourselves. Data processing details, dataset statistics, and training details are shown in Appendix A.
Baselines
We use multiple LLMs, including Qwen-Max [^2] https://dashscope.aliyuncs.com/compatible-mode/v1 Bai et al. (2023), ChatGPT, and GPT4 in all of our experiments, due to their strong reasoning ability. The used search tools include Quark, Bing, and Google. Considering that the LLM can answer easy queries independently, we set a non-retrieval baseline, i.e., the LLM directly responds to the query. Consequently, its usage cost is set to 0. In addition, since existing methods Shnitzer et al. (2023); Šakota et al. (2024) assign queries to the most suitable LLM, we devise "LLMs-UB", Specifically, For a query and corresponding scores obtained from several non-retrieval methods, such as Qwen-only, ChatGPT-only, and GPT4-only, we always assign the query to the LLM with the highest score. Therefore, this can be seen as an upper bound of Shnitzer et al. (2023); Šakota et al. (2024).
@@ 4.2.1 Metrics
We also report the average cost of each method. For simplicity, we set the average costs of Quark, Bing, and Google to 0.33, 2, and 1, respectively.
Methods PrivateTimeQA CDQA WebQA
Acc. Cost Acc. Cost Acc. Cost
ChatGPT
LLM-only 30.47 0 29.91 0 59.25 0
LLMs-UB 55.64 0 48.32 0 93.00 0
LLM+Quark 67.81 0.33 55.91 0.33 93.47 0.33
LLM+Bing 66.63 2 60.08 2 91.18 2
LLM+Google 68.63 1 61.13 1 94.07 1
Ours (Roberta-large) 68.89 1.24 63.28 0.68 93.55 0.52
GPT4
LLM-only 29.9 0 28.46 0 71.77 0
LLMs-UB 55.64 0 48.32 0 93.00 0
LLM+Quark 71.87 0.33 57.22 0.33 94.42 0.33
LLM+Bing 70.66 2 61.22 2 93.36 2
LLM+Google 73.35 1 64.42 1 95.5 1
Ours (Roberta-large) 73.6 1.01 65.74 0.58 94.54 0.79
Qwen-max
LLM-only 50.30 0 45.90 0 91.14 0
LLMs-UB 55.64 0 48.32 0 93.00 0
LLM+Quark 77.54 0.33 59.84 0.33 96.33 0.33
LLM+Bing 76.91 2 63.86 2 93.48 2
LLM+Google 77.95 1 66.75 1 94.97 1
Ours (Roberta-large) 78.41 1.23 69.56 1.20 94.68 1.30
Table 1: The result of our method when we use ChatGPT, GPT4 and Qwen-max for response generation. "LLM-only" means the non-retrieval baseline. "LLM+Quark", "LLM+Bing" and "LLM+Google" means the LLM always calls Quark, Bing, or Google to solve queries. “Acc. (%)" denotes accuracy. Scores with bold denote the best values among baselines.
Methods Model Size TimeQA CDQA
Acc. Cost Acc. Cost
ChatGPT
Ours (Roberta-base) 125M 68.68 1.24 62.3 0.68
Ours (Roberta-large) 355M 68.89 1.24 63.28 0.68
Ours (Qwen-0.5B) 0.5B 67.99 1.25 62.14 1.38
Ours (Qwen-1.8B) 1.8B 68.73 1.02 61.37 1.03
GPT4
Ours (Roberta-base) 125M 72.48 1.45 65.57 0.83
Ours (Roberta-large) 355M 73.6 1.01 65.74 0.58
Ours (Qwen-0.5B) 0.5B 72.11 1.32 63.16 1.42
Ours (Qwen-1.8B) 1.8B 73.06 1.10 64.76 1.07
Table 2: The result of different backbones.
@@ 4.2.2 Results
The result is shown in Table 1. Whether using Qwen-max, ChatGPT, or GPT4 for response generation, LLM+Google usually outperforms LLM+Bing, LLM+Quark, and LLM-only, which coincides with human intuition that Google is currently the best search tool in the world. RAG outperforms LLM-only by a large margin, proving the necessity of utilizing tools for response generation. In addition, we have the following observations.
(1) On PrivateTimeQA and CDQA, our method significantly outperforms the best baseline, i.e., "LLM+Google". This trend is observed across multiple LLMs. This is because despite "LLM+Google" achieving the highest overall accuracy, it still fails to solve some cases that "LLM+Quark" or "LLM+Bing" can handle. Specifically, the proportion of such samples is 6.6% in the CDQA testset. Our method can adaptively assign these queries to the optimal tools, leading to better overall results. (2) Compared with "LLM+Bing", our method achieves a better result with a lower cost. Compared with "LLM+Google", our method has a better performance with a approaching cost. This demonstrates the generality of our method for different search engines and LLMs.
@@ 4.3 Evaluation of the Adaptively Assignment
According to different strategies in Section 3.3, we assign test queries to the corresponding tools and calculate the average accuracy and usage cost [^3] Bing Search charges \$10 for 1000 requests, and Google Search charges \$5 for 1000 requests. . Cost–accuracy curves are plotted, which present the relationship between the average accuracy and the average cost per query needed.
@@
We plot cost-accuracy curves on the test sets of PrivateTimeQA and CDQA in Figure 3. "LLM+Google" has a higher accuracy and a lower cost compared to "LLM+Bing". We additionally have the following observations. (1) Based on our $\mathcal{M}$, "Best-Performance" achieves a higher accuracy than "LLM+Google", albeit at a higher cost. The accuracy increment comes from the supplement from "LLM+Bing", with a side-effect of higher cost. (2) Compared to both "LLM+Google" and "Best-Performance", the "Cost-Saving" strategy is able to achieve higher accuracy with a lower cost. This shows that "Cost-Saving" is more flexible and can effectively save costs.
(a) The curve on the CDQA testset.
(b) The curve on the PrivateTimeQA testset.
Figure 3: The cost-accuracy curves on PrivateTimeQA and CDQA when using Qwen-max.
@@ 4.4 Further Discussion
Impact of Different Backbones
We investigate the impact of different backbones for training our router. We consider Roberta-base(125M), Roberta-large(355M), Qwen-0.5B (0.5B), Qwen-1.8B(1.8B) for experiments. When using Roberta-large as the backbone, the result is better than that based on Roberta-base. But when using the larger Qwen-0.5B (0.5B), Qwen-1.8B (1.8B), even if we have tried various training techniques, the results do not increase significantly. We speculate that the decoder-only models are not suitable for this routing task which requires accurate numerical predictions. In addition, the larger model leads to higher latency. Therefore, we finally chose Roberta-large as our backbone. Due to limited computing resources, we are unable to attempt larger models such as Qwen-7b, Llama7b, etc.
@@ 5 Conclusion
Due to the high cost of search tools, our training data only contains approximately 16000 examples,. In the future, with a sufficient budget, we will conduct testing on more datasets. We use small encoding models to train our predictive model instead of LLMs, mainly for the following two reasons: (1) Tool selection has a high requirement for latency, while the speed of large models is relatively slow and does not meet the requirement. (2) Tool selection requires the model to predict the score of each tool solving each problem, which is actually a numerical prediction problem. However, the auto-regressive LLMs have significant issues in numerical prediction. In our experiment, the performance of LLMs was inferior to that of small encoding models.
Appendix A Data Processing Details and Statistics
Construction Details about PrivateTimeQA
PrivateTimeQA primarily consists of questions related to timeliness. We collected question-answer examples from public QA websites. We prioritize privacy protection and data integrity, ensuring that the dataset contains only open-domain QA content and excludes any harmful, or unethical material. PrivateTimeQA is used to evaluate LLMs’ capability to answer timeliness questions. Due to confidentiality requirements, we are unable to publicly disclose the dataset.